A positive integer is twice another. The difference of the reciprocals of the two positive integers is 1/10. Find the two integers.
Answer by Boreal(15235)

numbers are x and 2x
the reciprocals are (1/x) and (1/2x)
(1/x)-(1/2x)=0.1
multiply everything by 2x
2-1=.2x
x=5
numbers are 5, 10
(1/5)-(1/10)=0.1